# Integrating Agent-Based and Compartmental Models for Infectious Disease Modeling: A Novel Hybrid Approach

### v1

@@ Abstract
This study investigates the spatial integration of agent-based models (ABMs) and compartmental models in infectious disease modeling, presenting a novel hybrid approach and studying its implications. ABMs, characterized by individual agent interactions and decision-making, offer detailed insights but are computationally intensive for large populations. Compartmental models, based on differential equations, provide population-level dynamics but lack granular detail. Our hybrid model aims to balance the granularity of ABMs with the computational efficiency of compartmental models, offering a more nuanced understanding of disease spread in diverse scenarios, including large populations.
We developed a custom ABM and a compartmental model, analyzing their infectious disease dynamics separately before integrating them into a hybrid model. This integration involved spatial coupling of discrete and continuous populations and evaluating the consistency of disease dynamics at the macro scale. Our key objectives were to assess the effect of model hybridization on resulting infection dynamics, and to quantify computational cost savings of the hybrid approach over the ABM. We show that the hybrid approach can significantly reduce computational costs, but is sensitive to between-model differences, highlighting that model equivalence is a crucial component of hybrid modeling approaches. The code is available at http://github.com/iebos/hybrid_model1.
@@ 1 Introduction
Preventing the widespread transmission of infectious diseases is a critical objective for health systems worldwide. Preemptively addressing and mitigating their spread is crucial for safeguarding public health, and disease modeling plays an essential role in this effort. By analyzing the societal impacts and developing effective intervention strategies, disease modeling has become a necessity, as underscored by the COVID-19 pandemic [1]. Tools, such as agent-based models (ABMs) and compartmental models, are fundamental approaches for understanding and managing disease spread [2]. As a result, these models have gained significant recognition in epidemiology and are invaluable to policymakers and public health officials [3].
Compartmental models are often used for mathematical modeling of infectious diseases [4] and provide insights into the spreading dynamics on a macro-scale. They are often based on (deterministic) ordinary differential equations (ODEs) and describe the rates of movement between different compartments, such as susceptible, infected, and recovered, within a population. This modeling approach offers valuable perspectives on overall population-level dynamics [5, 6, 7]. Although based on the simplistic assumption of a homogeneously mixed population, their complexity and spatial resolution can be increased by including human mobility in meta-population models [8, 9, 10] or by coupling populations on multiple layers [11]. In contrast, ABMs simulate the behavior of individual agents, accounting for their interactions, movements, and decision-making processes. This enables a detailed analysis of disease transmission dynamics under various scenarios and provides a more granular and realistic approach to infectious disease modeling [12, 13]. This allows for population heterogeneity, for example, by representing the population as a network of agents [14], or by capturing complex spatial-temporal patterns through explicit modeling of movement and interactions [15, 16, 17].
However, ABMs require significantly more computational resources compared to (ODE-based) compartmental models, a discrepancy that grows with population size [for a comparison, see 18, 19]. Further, ABMs often require more complex, fine-grained data for fitting or verification [2]. This hinders ABM-based analysis of large populations and calls for new strategies that combine the computational efficiency of ODE-based approaches with the detailed insights of ABM-based approaches. Such strategies could be particularly valuable to policymakers during health emergencies such as the Covid-19 pandemic.
@@
In this paper, we propose a hybrid modeling approach that spatially integrates an ABM with an ODE-based compartmental model. Such a hybrid model could be utilized to integrate multiple geographic regions with different population characteristics and data availability, while allowing for population movement between regions. This hybrid approach could reduce the computational costs associated with a pure ABM while retaining analytical insights. Moreover, it may scale up to larger populations without sacrificing significant predictive power, as would be the case with a purely compartmental model. To facilitate such applications, we systematically explore the proposed hybrid model and evaluate its feasibility.
Despite their potential, hybrid models have received relatively little attention in infectious disease modeling research. Few studies have explored hybrid modeling approaches, most of which have not critically examined how model coupling impacts the dynamics of infectious diseases. One of the main goals of this paper is to analyse coupling effects on infectious disease dynamics and the computational efficiencies of the hybrid model compared to a pure ABM approach. For this, we study two scenarios in a simplified toy example: In the first, movement within the ABM component is random, resulting in a homogeneous population. In the second, agent-movement is influenced by a landscape that encourages spatial heterogeneity, and therefore deviates from the ODE-component.
One of the primary strengths of this study is its methodical assessment of the effects of coupling on key indicators that reflect the infectious disease dynamics within a spatially coupled hybrid model. We find that the spatial location of the coupling mechanism affects resulting infectious disease dynamics when agent movement is non-random. This finding underscores the importance of understanding the effects of model coupling for accurate epidemic modeling.
This study addresses the following questions: (1) How does spatial coupling of an ABM with an ODE-based model affect the dynamics of infection spread in the overall population? (2) How much computational cost is saved with the hybrid model compared to a pure ABM?
@@
The paper is structured as follows: We begin by reviewing related work on hybrid models that couple compartmental and agent-based models, particularly in the field of infectious disease modeling. Next, we detail the implementation of the ABM, the compartmental model, and the hybrid model, with an explicit description of the coupling mechanism. Following, we vary the weight of the compartmental component within the hybrid model in two experiments to assess the error and infection dynamics of the hybrid model. The first experiment deals with a homogeneous population scenario, while the second experiment deals with a hybrid model with a spatially heterogeneous ABM-component. Finally, we discuss the results and give an outlook for future research.
@@ 2 Related Work
One of the earliest hybrid model approach for infectious diseases is described by Bobashev et al. [20]. In that study, a temporal coupling was implemented, meaning that the model alternates between an ABM state and a compartmental model state, with the switch being triggered when the number of infected individuals crosses a threshold value. Aggregate population values of the ABM are used as initial conditions for the compartmental model state. The authors argued that insights into epidemic processes at the local level diminish in informative value with an increasing number of infected individuals. They evaluated several threshold values and demonstrated how such a model can be applied to a single hypothetical city. Further, using air travel data, they extended this model to a network of cities. [21] followed a similar approach, applying the switching method to a meta-population setting for several regions in Ireland. They compared computational execution times and highlighted the potential of this model to conserve resources. However, they also noted that the computationally most efficient configurations might yield inferior model results due to the suboptimal performance of the compartmental model component, measured by the total number of infected individuals and the length of the outbreak.
Rather than using infection numbers, Nghi et al. [22] implemented a hybrid model where the state switch is triggered when local agent density crosses a threshold. The authors argued that dense clusters of agents can be viewed as a homogeneous population, and therefore a compartmental model is sufficient.
@@
Another approach, found in Bradhurst et al. [23] and Banos et al. [24], involves using agents to couple compartmental models in a hybrid model. In Bradhurst et al. [23], agents represent herds of livestock, and within each agent, ordinary differential equations (ODEs) determine the infection metrics over time. Agent-agent interactions govern the spread of diseases between herds through different pathways. In Banos et al. [24], cities are represented by nodes in a network. The infectious disease dynamics within the cities are modeled with a compartmental SIR-model. A meta-population approach was compared to a hybrid model, where the nodes are connected through agent-based air travel: Agents can fly from one city to another, and infection events are possible within the aircraft. The spread of a disease can be slower in the hybrid approach if only one or a few infected agents are localized on one node. This is because diffusion happens evenly in the meta-population model with constant mobility rates for all compartments, while mobility is discrete and stochastic in the hybrid approach. A similar coupling mechanism was employed in Marilleau et al. [25], focusing on modeling the spread of a vole population in a French region. This spatial model divides the region into grid cells, allowing vole agents to travel between those cells. Within each cell, the vole population size is determined by equation-based models. Once the population of a cell reaches a critical value, the cell generates vole agents that leave the cell and migrate to another cell with more suitable conditions.
@@
A problem that arises when coupling an ABM with a compartmental model is the differing population shape in each model type. In an ABM, the population is represented by distinct entities, while an ODE model represents the population as a continuous matter [26]. Sewall et al. [27] highlight that this is particularly a problem when agents travel out of a compartmentally modeled region due to information loss when converting agents to a continuous population. Addressing this disparity is crucial in developing a successful hybrid model, as it involves accurately transitioning the population between discrete entities and continuous mass representations.
@@
In conclusion, while various hybrid models have been developed to leverage the strengths of both ABMs and compartmental models, they often face challenges such as determining appropriate switching mechanisms and managing the transition between discrete and continuous representations of the population. In the following section, we will describe our suggested hybrid model and the approaches we use to address these challenges effectively.
@@ 3 The Hybrid Model
This section introduces our proposed hybrid modeling approach, which combines the strengths of a microscopic ABM with a macroscopic compartmental ODE model. This integration aims to achieve a balance between computational efficiency and detailed, realistic representation of infection dynamics. In the hybrid model, infectious disease dynamics are modeled by two components in parallel: In the first component, infections spread through agent-agent interactions in a continuous coordinate space. The second component employs ODEs to model these dynamics. This approach leverages the fine-grained interactions and spatial detail of ABMs while utilizing the computational simplicity of compartmental models to scale up to larger populations effectively.
In the following sections, we provide a detailed description of the agents, the environment, and the coupling mechanism to illustrate how our hybrid model addresses the challenges of combining discrete and continuous population representations. Moreover, we demonstrate the model’s potential to reduce computational costs while maintaining accurate and insightful predictions of infectious disease dynamics.
@@ 3.1 Sub-Model M1: Agent-Based Model (ABM)
The ABM simulates agents within a 2D rectangular region in a xy-plane, defined by the coordinates $0\leq x\leq 9$ and $0\leq y\leq 9$. Each agent occupies a $(x,y)$-coordinate within this space and is in one of three states: susceptible, infectious, or recovered. The probability $p_{i}$ of a susceptible agent $i$ to become infectious within a time step is $p_{i}=1-(1-\rho)^{k}$, where $\rho$ is the base probability and $k$ is the number of infectious agents within a distance of $\theta$ to the susceptible agent. This transmission model assumes a fixed transmission probability upon contact and increases with repeated exposure, akin to a contact-duration model [see 28, 29]. While dose-response models are more realistic [30, 31], we chose to implement this simplified model due to its similarity to the transmission mechanism in ODE models: In the ODE model, the transmission rate is a function of the contact rate and the infection probability. In this version of an ABM, the contact rate can be seen as a function of the radius size and the step size. Infectious agents recover at a rate of $1/D$ per time step, where $D$ is the average time an agent remains in the infectious state, becoming immune to subsequent infections through agent-agent interactions.
@@ 3.1.1 Agent movement
In variant M1-B, however, we implemented a landscape that influences agent movement [see 32, 33]. This landscape simulates non-random movement and is modeled by a bivariate density function
$f(x,y)=-e^{\frac{\left(-(x-X)^{2}+(y-Y)^{2}\right)}{\eta}}$ (1)
@@
under periodic boundary conditions. Agents switch to the repulsion state once they approached the center of the valley within a distance of 0.3 and switch back to attraction at a rate of $1/(3\eta)$. This dynamic switching mechanism allows for movement across the simulation space, resembling commute-like behavior, instead of the entire population remaining around the center for the entire simulation time. A snapshot of a realization of the ABM at $t=9$ can be found in Figure 2, showing a dense cluster of agents at the center, surrounded by less densely populated space.
Figure 2: Snapshot ($t=9$) of a realization ABM variant B. Each point represents the position of a susceptible (yellow), infectious (red), or recovered (blue) agent.
@@ 3.1.2 Initialization
For initialization, agents are placed randomly across the simulation space in variant A. In variant B, initial positions and behavior states are determined using the Monte Carlo method, based on the bivariate density function for the coordinates, and $1/(1-1/3\eta)$ for the behavior state probabilities, and the outbreak starts after a burn-in period.
@@ 3.2 Sub-Model M2: Compartmental Model
The compartmental model follows the classic SIR (Susceptible-Infectious-Recovered) framework, describing the dynamics of disease spread in a population with ordinary differential equations (ODEs) [34, 5]. The model divides the population into three compartments: susceptible individuals ($S$), infectious individuals ($I$), and recovered individuals ($R$). The dynamics of the SIR model are governed by a system of ordinary differential equations (ODEs). The rate of change for each compartment is given by:
@@
where $\beta$ represents the transmission rate of the disease, $\gamma$ is the recovery rate, and $N$ is the total population size. This model captures the rates at which the population transitions between compartments. Given initial conditions, step size, and time frame, the ODEs can be solved numerically, providing compartment sizes over time - and thus simulating the dynamics of disease spread in a population.
@@ 3.3 Coupling
In the hybrid model, the population-space is spatially divided into an ABM (M1) and a compartmental region (M2), while allowing for population flow between regions. The interface between M1 and M2 is a vertical line that can be moved along the x-axis to adjust the proportions (or weight) of each model component. This flexible interface allows for varying the influence of each model within the hybrid framework.
The temporal resolution and the order of progresses can significantly impact the results of both ABMs and compartmental models [18]. This aspect is particularly crucial in our hybrid model due to the interaction between a discrete model and a continuous model [26]. To maintain coherence, time is implemented as a discrete variable in both models, and interactions within the ABM and between models occur synchronously. Figure 3 visualizes the coupling mechanism in the hybrid model, depicting the sequence of processes within a single time step.
The subsequent sections will detail these processes.

### v2

@@ Abstract
This study investigates the spatial integration of agent-based models (ABMs) and compartmental models for infectious disease modeling, presenting a novel hybrid approach and examining its implications. ABMs offer detailed insights by simulating interactions and decisions among individuals but are computationally expensive for large populations. Compartmental models capture population-level dynamics more efficiently but lack granular detail. We developed a hybrid model that aims to balance the granularity of ABMs with the computational efficiency of compartmental models, offering a more nuanced understanding of disease spread in diverse scenarios, including large populations. This model spatially couples discrete and continuous populations by integrating an ordinary differential equation model with a spatially explicit ABM. Our key objectives were to systematically assess the consistency of disease dynamics and the computational efficiency across various configurations. For this, we evaluated two experimental scenarios and varied the influence of each sub-model via spatial distribution. In the first, the ABM component modeled a homogeneous population; in the second, it simulated a heterogeneous population with landscape-driven movement. Results show that the hybrid model can significantly reduce computational costs but is sensitive to between-model differences, highlighting the importance of model equivalence in hybrid approaches. The code is available at: git.zib.de/ibostanc/hybrid_abm_ode
@@ 1 Introduction
Preventing or mitigating the widespread transmission of infectious diseases is a critical objective for health systems worldwide. Disease modeling is essential to develop effective intervention strategies and analyze their societal impacts, as underscored by the COVID-19 pandemic [1]. In this context, agent-based models (ABMs) and compartmental models are fundamental approaches [2]. As a result, these models have gained significant recognition in epidemiology and are invaluable to policymakers and public health officials [3].
Compartmental models are traditionally used for mathematical modeling of infectious diseases and capture high-level system dynamics by describing the rates of transition between sub-populations, such as susceptible, infected, and recovered [4]. They offer valuable insights on spreading dynamics at the macro-scale, but their reliance on a homogeneously mixed population limits realism [5, 6, 7]. Although meta-population models can increase their complexity and spatial resolution by further disaggregating the population (e.g. through mobility [8, 9, 10] or layered populations [11]), they retain the simplifying assumption of homogeneity within each sub-population.
In contrast, ABMs simulate the behavior of individual agents, accounting for their interactions, movements, and decisions. This enables a detailed analysis of disease transmission under various scenarios, providing a more granular and realistic approach to infectious disease modeling [12, 13]. ABMs allow for population heterogeneity, for example, by representing the population as a network of agents [14], or by capturing complex spatial-temporal patterns through explicit modeling of movement and interactions [15, 16, 17]. This makes ABMs valuable for exploring individual-level variability and allows for a straightforward implementation of policy interventions.
However, this increased realism comes at a cost: ABMs often require more complex, fine-grained data for fitting and verification [2]. While their emergent macro-level outcomes can be validated with observed data, their implemented micro-level relationships are often based on unobserved assumptions [18, 19], making them less adept to derive underlying parameters than compartmental models, such as the basic reproduction number [20]. Further, ABMs require significantly more computational resources compared to compartmental models, a discrepancy that grows with population size [21, 22].
@@
These challenges hinder ABM-based analysis of large populations and call for a hybrid modeling approach that combines the computational and analytical efficiency of compartmental models with the detailed insights of ABMs. Such a hybrid model could integrate multiple geographic regions with different population characteristics and data availability, while allowing for population movement between regions. Moreover, it may scale up to larger populations without sacrificing significant predictive power.
Despite their potential, hybrid models have received relatively little attention in infectious disease modeling research. While the meta-population approach couples sub-populations under the same modeling paradigm and retains the assumption of homogeneity within these, hybrid approaches could flexibly loosen or constrain assumptions for different populations where appropriate, harnessing the strengths and data requirements of both model types. Few studies have explored such hybrid models, most of which have not critically examined how model coupling impacts the dynamics of infectious diseases under varying conditions.
To fill this gap, we propose a hybrid model that spatially integrates an ABM with a compartmental model based on ordinary differential equations (ODEs). One of the primary strengths of this study is its methodical assessment of the effects of spatial coupling on key indicators that reflect the infectious disease dynamics. Further, we specifically address challenges in coupling discrete and continuous population representations under different modeling paradigms and systematically explore the feasibility of the hybrid approach by addressing the following: (1) How does spatial coupling of an ABM with an ODE-based model affect infection spread dynamics? (2) What are the computational cost savings compared to a pure ABM?
To answer these, we study two scenarios in a simplified toy example: In the first, random movement in the ABM component results in a homogeneous population. In the second, landscape-influenced movement encourages spatial heterogeneity, and therefore deviates from the ODE component. In both scenarios, we systematically alter the spatial distribution and the interface location between sub-models and examine the resulting impact on key indicators for infection dynamics and computational efficiency. Our results show that the spatial location of the coupling mechanism can affect the peak time and shape of the infectious curve, particularly when movement patterns differ across regions. These findings underscore the importance of understanding the effects of hybrid model coupling for accurate epidemic modeling.
@@
The article is structured as follows: We begin by reviewing related work on hybrid models that couple compartmental models with ABMs, particularly in the field of infectious disease modeling. Next, we detail the implementation of the ABM, the compartmental model, and the hybrid model, with an explicit description of the coupling mechanism. Following, we vary the spatial distribution of model components within the hybrid model in two experiments to assess the consistency of infection dynamics across configurations. Finally, we discuss the results and give an outlook for future research.
@@ 2 Related Work
One of the earliest hybrid model approaches for infectious diseases is described by Bobashev et al. [23]. This study implemented a temporal coupling that alternates the model between an ABM state and a compartmental model state, with the switch being triggered when the number of infected individuals crosses a threshold value. Aggregate population values of the ABM were used as initial conditions for the compartmental model state. The authors argued that insights into epidemic processes at the local level diminish in informative value with an increasing number of infected individuals. They evaluated several threshold values and demonstrated how such a model can be applied to a single hypothetical city. Further, using air travel data, they extended this model to a network of cities. [24] followed a similar approach, applying the switching method to a meta-population setting for several regions in Ireland. They compared computational execution times and highlighted the potential of this model to conserve resources. However, they also noted that the computationally most efficient configurations might yield inferior model results due to the suboptimal performance of the compartmental model component, measured by the length and final size of the outbreak. Rather than using infection numbers, Nghi et al. [25] implemented a hybrid model that triggers a state switch when local agent density crosses a threshold. The authors argued that dense clusters of agents can be viewed as a homogeneous population, and therefore a compartmental model is sufficient.
@@
Another approach involves using agents to couple compartmental models in a hybrid model. For example, Nguyen et al. [26] investigated the role of temporary staff during the COVID-19 pandemic by modeling care homes as compartmental models and temporary staff as agents moving between them. Similarly, in Bradhurst et al. [27], agents represent herds of livestock, and within each agent, ODEs determine infection dynamics. Agent-agent interactions govern the spread of diseases between herds through different pathways. In Banos et al. [28], cities are represented as network nodes with infectious disease dynamics modeled by compartmental models. A hybrid model, where agent-based air travel connects nodes with possible transmission during flight, was compared to a meta-population approach. The hybrid model showed slower infection spread for scenarios with few infected agents per node due to the discrete, stochastic mobility implementation, contrasting the even diffusion in the deterministic meta-population model. A similar coupling mechanism was employed in Marilleau et al. [29] to model the spread of a vole population in a French region. This spatial model divides the region into grid cells, allowing vole agents to travel between those cells. Within each cell, the vole population size is determined by equation-based models. Once the population of a cell reaches a critical value, the cell generates vole agents that leave the cell and migrate to another cell with more suitable conditions.
@@
A problem that arises when coupling an ABM with a compartmental model is the differing population shape in each model type. In an ABM, the population is represented by distinct entities, while an ODE model represents the population as continuous matter [30]. Sewall et al. [31] highlight that this is particularly problematic when agents leave a compartmentally modeled region due to information loss when converting agents to a continuous population. In Solano et al. [32], this caused a total population decline because the agent-population size was fixed while compartmental population sizes were unconstrained, leading to asymmetric population flow. Additionally, individuals entering the ABM from a compartmental node were represented in aggregated floating point numbers, and reconstructing these into integer agent counts lead to discrepancies. Addressing this disparity is crucial in developing a successful hybrid model, as it involves accurately transitioning the population between discrete entities and continuous mass representations.
In a review of hybrid models in operational research, Brailsford et al. [18] note that published studies often have limited reporting on verification and validation, with many failing to sufficiently describe the architecture of the hybrid models used. Although their review focused on operational research, we found similar issues in the hybrid infectious disease models reviewed in this section. Specifically, these models frequently omit details on interactions between components and the mechanisms of data exchange between sub-models. Additionally, as with the studies in Brailsford et al. [18], the implications of these hybrid models for infectious disease prediction and control are often underexplored, a notable gap given the importance of model accuracy in public health contexts.
@@
In conclusion, while various hybrid models have been developed to leverage the strengths of both ABMs and compartmental models, they often face challenges such as determining appropriate switching mechanisms and managing the transition between discrete and continuous representations of the population. In the following section, we describe our proposed hybrid model and how we addressed these challenges effectively.
@@ 3 The Hybrid Model
This section introduces our proposed hybrid modeling approach, which combines the strengths of a microscopic ABM with a macroscopic compartmental ODE model. The hybrid approach aims to balance computational efficiency with a detailed and realistic representation of infection dynamics. In the hybrid model, infectious disease dynamics are modeled by two parallel components: the first simulates infections via agent-agent interactions in a continuous coordinate space; the second employs ODEs to model these dynamics. This design leverages the fine-grained interactions and spatial detail of ABMs while utilizing the computational simplicity of compartmental models to scale up to larger populations effectively.
The following sections detail the agents, environment, and coupling mechanism, illustrating how our hybrid model addresses the challenges of combining discrete and continuous population representations. Moreover, we demonstrate the model’s potential to reduce computational costs while maintaining accurate and insightful predictions of infectious disease dynamics.
@@ 3.1 Sub-Model M1: Agent-Based Model (ABM)
The ABM simulates agents within a 2D rectangular region in a xy-plane, defined by the coordinates $0\leq x\leq 9$ and $0\leq y\leq 9$. Each agent occupies a $(x,y)$-coordinate within this space and is in one of three states: susceptible, infectious, or recovered. The probability $p_{i}$ of a susceptible agent $i$ to become infectious within a time step is $p_{i}=1-(1-\rho)^{k}$, where $\rho$ is the base probability and $k$ is the number of infectious agents within a distance of $\theta$ to the susceptible agent. This transmission model assumes a fixed transmission probability upon contact and increases with repeated exposure, akin to a contact-duration model [see 19, 33]. While dose-response models are more realistic [34, 35], we chose to implement this simplified model due to its similarity to the ODE’s transmission mechanism: In the ODE model, the transmission rate is a function of the contact rate and the infection probability. In our ABM, the contact rate can be seen as a function of the radius size and the step size. Infectious agents recover at a rate of $1/D$ per time step, where $D$ is the average time an agent remains in the infectious state, becoming immune to subsequent infections through agent-agent interactions.
@@ 3.1.1 Agent Movement
In variant M1-B, however, we implemented landscape-driven agent movement [see 36, 37]. This landscape induces non-random movement and is generated by a bivariate density function
$f(x,y)=-e^{-\frac{(x-X)^{2}+(y-Y)^{2}}{\eta}}$ (1)
@@
under periodic boundary conditions. Agents switch to the repulsion state once they approached the center of the valley within a distance of 0.3 and switch back to attraction at a rate of $1/(3\eta)$. This dynamic switching mechanism allows for movement across the simulation space, resembling commute-like behavior, instead of the entire population remaining around the center for the entire simulation time. A snapshot of an M1-B realization at $t=9$ can be found in Figure 2, showing a dense cluster of agents at the center, surrounded by less densely populated space.
Figure 2: Snapshot of an ABM realization (variant M1-B) at $t=9$. Each point represents the position of a susceptible (yellow), infectious (red), or recovered (blue) agent.
@@ 3.1.2 Initialization
For initialization, agents are placed randomly across the simulation space in M1-A. In M1-B, initial positions and behavior states are determined using the Monte Carlo method, based on the bivariate density function for the coordinates and $1/(1-1/3\eta)$ for the behavior state probabilities, with the outbreak starting after a burn-in period.
@@ 3.2 Sub-Model M2: Compartmental Model
The compartmental model follows the classic SIR (Susceptible-Infectious-Recovered) framework, describing the dynamics of disease spread in a population with ODEs [38, 5]. The model divides the population into three compartments: susceptible individuals ($S$), infectious individuals ($I$), and recovered individuals ($R$). The rates at which the population transitions between compartments are given by:
@@
where $\beta$ represents the transmission rate of the disease, $\gamma$ is the recovery rate, and $N$ is the total population size. Given initial conditions, step size, and time frame, the ODEs can be solved numerically, providing compartment sizes over time - and thus simulating the dynamics of disease spread in a population.
@@ 3.3 Coupling
In the hybrid model, the population-space is spatially divided into an ABM (M1) and a compartmental region (M2), while allowing for population flow between regions. The interface between M1 and M2 is a vertical line that can be moved along the x-axis to adjust the proportions (or weight) of each sub-model. This flexible interface allows us to vary the influence of each model within the hybrid framework.
The temporal resolution and the order of progresses can significantly impact the results of both ABMs and compartmental models [21]. This aspect is particularly crucial in our hybrid model due to the interaction between a discrete model and a continuous model [30]. As we implemented discrete time steps in the ABM, we maintain coherence by iteratively solving the ODE model for one time step while interactions within the ABM and between sub-models occur synchronously.
As discussed earlier, the exchange between discrete and continuous population representations requires careful handling to ensure accurate population tracking and infection dynamics. Figure 3 visualizes the coupling mechanism in the hybrid model, depicting the sequence of processes within a single time step. The subsequent sections detail these processes.